Derive and solve the equation of motion of a particle, in a uniform gravita-

tional field, projected with initial horizontal velocity v 0 at a height h.


1
Expert's answer
2021-08-16T08:37:12-0400

"v_x=v_0\cos \alpha,"

"v_y=v_0\sin \alpha-gt,"


"x=v_0\cos\alpha t,"

"y=y_0+v_0\sin\alpha t-\frac{gt^2}2,"


"y(x)=y_0+\tan\alpha x-\frac{gx^2}{2v_0^2\cos^2\alpha},"


"\alpha=0,"

"y_0=h,"


"y(x)=h-\frac{gx^2}{2v_0^2}."